10.48 From her firm’s computer telephone log, an executive found that the mean length of 64 telephone calls during July was 4.48 minutes with a standard deviation of 5.87 minutes. She vowed to make an effort to reduce the length of calls. The August phone log showed 48 telephone calls whose mean was 2.396 minutes with a standard deviation of 2.018 minutes.
(a) State the hypotheses for a right-tailed test.
(b) Obtain a test statistic and p-value assuming unequal variances. Interpret these results using α = .01.
(c) Why might the sample data not follow a normal, bell-shaped curve? If not, how might this affect your conclusions?
Thank you, Stan

Answer by stanbon(48568)

From her firm’s computer telephone log, an executive found that the mean length of 64 telephone calls during July was 4.48 minutes with a standard deviation of 5.87 minutes. She vowed to make an effort to reduce the length of calls. The August phone log showed 48 telephone calls whose mean was 2.396 minutes with a standard deviation of 2.018 minutes.
(a) State the hypotheses for a right-tailed test.
-------------------------
Note: She wants to reduce the call length so the test should be left-tailed
------------------------------------------------------
Ho: u = 4.48
Ha: u < 4.48
-----------------

(b) Obtain a test statistic and p-value assuming unequal variances. Interpret these results using α = .01.
TS: z(2.396) = (2.396-4.48)/[2.018/sqrt(48)] = -7.1548
p-value = P(z < -7.1548) = 0.0000000..
-----------------------

(c) Why might the sample data not follow a normal, bell-shaped curve? If not, how might this affect your conclusions?
The calls may not be those of a very large population and therefore
lack the normal symmetry of distribution.
==================================================
Cheers,
Stan H.